# Constraining $\boldsymbol{|V_{cs}|}$ and physics beyond the Standard Model from exclusive (semi)leptonic charm decays

### v1

@@ 1 Introduction
The Cabibbo-Kobayashi-Maskawa (CKM) quark mixing matrix is a central parameter in the Standard Model of particle physics (SM). Precision determinations of its elements are key to understanding the origin of (quark) flavour. The CKM element $|V_{cs}|$ is currently the second-most precisely known CKM element. The Particle Data Group (PDG) has determined a reference value for $|V_{cs}|$ with a relative uncertainty of $0.6\%$ PDGCKMReview . In computing this value, contemporary world averaged of the total branching ratios of $D\to\bar{K}\ell^{+}\nu$ and $D_{s}^{+}\to\ell^{+}\nu$ decays are used in combination with theoretical determinations of the relevant hadronic form factors and decay constants.
In this work, we study exclusive flavour-changing $c\to s\ell^{+}\nu$ processes in a global fit. These processes include $\bar{D}_{s}\to\ell^{+}\nu$, $\bar{D}_{s}^{*}\to\ell^{+}\nu$, $D\to\bar{K}\ell^{+}\nu$, and $\Lambda_{c}\to\Lambda\ell^{+}\nu$. First, we determine if the available data can be simultaneously described by the available theory predictions. This question is not trivial, since the hadronic matrix elements are connected to each other within the framework of dispersive bounds Okubo:1973tj ; Boyd:1997kz ; Caprini:1997mu .
We determine $|V_{cs}|$ from the global fit within this setup and discuss the agreement of the result with CKM unitarity. We also probe if determinations of $|V_{cs}|$ from the individual decay processes are in agreement with each other and with the quoted PDG value PDGCKMReview . Finally, we determine how Beyond the Standard Model (BSM) effects are constrained by $c\to s\ell^{+}\nu$ processes by determining the maximally allowed parameter space within the $sc\nu\ell$ sector of the Weak Effective Theory (WET) Aebischer:2017gaw ; Jenkins:2017jig ; Jenkins:2017dyc . Our analysis leverages procedures and techniques first developed in the context of a study of $b\to u\ell^{-}\bar{\nu}$ decays Leljak:2023gna .
@@ 2.1 Theoretical Framework
where $\mathcal{C}^{\ell}_{i}$ are the Wilson coefficients and $\mathcal{O}_{i}$ the local field operators. The normalisation in terms of the Fermi constant $G_{\textrm{F}}$ and an arbitrary constant $\tilde{V}_{cs}$ proves to be convenient later on. The Wilson coefficients encode the dynamics of the full theory, either the SM or any viable BSM theory, above the separation scale $\mu_{c}\simeq 1.275\,\textrm{GeV}$. The matrix elements of the operators encode the dynamics below the separation scale. Here, we assume that there are no right-handed neutrinos of mass less than $\mu_{c}$. As a consequence, the basis of operators of mass dimension six reduces to five independent operators. Our choice of basis reads
@@
while all other Wilson coefficients are zero. In the SM, we identify $\tilde{V}_{cs}^{*}\mathcal{C}^{\ell}_{V,L}$ with the conjugated CKM matrix element $V_{cs}^{*}$. Beyond the SM, the Wilson coefficients provide a low-energy footprint of the genuine BSM dynamics at or above the electroweak scale. These coefficients can then be used to constrain the parameters of a UV-complete BSM model.
@@ 2.2 Approach
Here, $\vec{\vartheta}$ represents the parameters of interest and $\vec{\nu}$ represents the nuisance parameters. The latter arise exclusively from our description of hadronic matrix elements in the observables considered here. The posterior PDF is proportional to the likelihood function $P(D\,|\,\vec{\vartheta},\vec{\nu},M)$, which encapsulates the experimental and theoretical constraints imposed by the data $D$ under consideration, and to the prior PDF $P_{0}(\vec{\nu},\vec{\vartheta}\,|\,M)$, which accounts for our prior knowledge about the parameters in the model $M$. We discuss the various datasets and fit models entering our analysis in section 2.3 and section 2.4, respectively.
@@
Their ratio, the so-called Bayes factor $K\equiv P(D\,|\,M_{1})/P(D\,|\,M_{2})$, then provides information about the efficiency of the two models in describing the common dataset $D$. Following Jeffreys’ interpretation Jeffreys:1939xee , the model $M_{1}$ is preferred over the model $M_{2}$ if $K>1$. This preference can be characterised either as strong if $10010<K<$ 100 or as decisive if $100<K$. A Bayes factor $103<K<$ 10 is still interpreted to be substantially in favour of $M_{1}$ over $M_{2}$, whereas a Bayes factor of $31<K<$ 3 is “barely worth mentioning”. For $K<1$, the above interpretation holds in favour of $M_{2}$ over $M_{1}$ with the replacement $K\to 1/K$.
@@ 2.3 Experimental Data
However, the total decay width of the $D_{s}^{*+}$ is currently not known from any experiment. Instead, we use the prediction for the decay width determined in Ref. Meng:2024gpd . This value is obtained from the PDG world average of the experimental measurements of the dominant branching ratio $\mathcal{B}(D_{s}^{*+}\to D_{s}^{+}\gamma)$ ParticleDataGroup:2022pth and a lattice QCD prediction for the partial decay width $\Gamma(D_{s}^{*+}\to D_{s}^{+}\gamma)$ Meng:2024gpd .
@@
which is based on results by the Belle Belle:2006idb , BES BES:2004rav , BESIII BESIII:2015tql ; BESIII:2021mfl , and CLEO-c CLEO:2009svp experiments. This average dominated by the BESIII measurement in Ref. BESIII:2015tql .
@@
based on measurements from BES BES:2004obp , BESIII BESIII:2015jmz ; BESIII:2016hko ; BESIII:2017ylw ; BESIII:2021mfl . The differential decay rate is also available from Ref. BESIII:2017ylw , used in form of the normalised decay rate $1/\Gamma\,d\Gamma/dq^{2}$. For $D^{+}\to\bar{K}^{0}\mu^{+}\nu$, we use the only available results coming from BESIII BESIII:2016gbw ,
@@
For $\Lambda_{c}\to\Lambda\{e^{+},\mu^{+}\}\nu$ decays, we use measurements of the branching ratios by the BESIII experiment BESIII:2022ysa ; BESIII:2016ffj , yielding
@@
$\mathcal{B}(\Lambda_{c}\to\Lambda\mu^{+}\nu)$ $=(3.49\pm 0.46\pm 0.27)\%\,.$
Although measurements of the differential distributions for the decay chain $\Lambda_{c}\to\Lambda(\to p\pi)e^{+}\nu$ have been undertaken by BESIII BESIII:2022ysa , the data have not been made public. The differential data strongly constrain the BSM parameters space, as discussed in the context of $\Lambda_{b}\to\Lambda_{c}(\to\Lambda\pi)\ell^{-}\bar{\nu}$ Boer:2019zmp ; the analogue of the $\Lambda_{c}$ decays among semileptonic $b\to c\ell^{-}\bar{\nu}$ processes.
@@
We do not use decays to $\eta^{(\prime)}$ because their form factors, while available from light-cone sum rules and lattice QCD Duplancic:2015zna ; Hu:2021zmy ; Bali:2014pva , are not yet as stringently constrained as the ones for the other modes used in this analysis. We also do not use measurements of $D^{+}\to\bar{K}^{*0}\ell^{+}\nu$ nor $D_{s}^{+}\to\phi\ell^{+}\nu$, because their hadronic final states are unstable vector resonances. As such, they may suffer from $S$-wave pollution, which is a problem from the point of view of both the experimental extraction of the observables and the theoretical predictions. Therefore, we consider them not to be competitive with processes included above.
@@ 2.4 Parameters of Interest and their Priors
This model has no parameters of interest. We use a fixed normalisation $\tilde{V}_{cs}=0.975$, corresponding to the average determined by the Particle Data Group PDGCKMReview . The left-handed Wilson coefficient $\mathcal{C}^{\ell}_{V,L}$ is set to its SM value in (3), and all other Wilson coefficients are fixed to zero. For the purpose of our model comparisons, this fit model serves as the null hypothesis.
@@
This model has nine parameters of interest, which describe the degrees of freedom for the five complex Wilson coefficients that appear in the effective Hamiltonian eq. 1. The tenth degree of freedom, which can be chosen by imposing $\arg\tilde{V}_{cs}=0=\arg{\mathcal{C}^{\ell}_{V,L}}$, is fixed, since the overall phase of the effective Hamiltonian is unobservable. To provide an absolute scale for the magnitude of these Wilson coefficients, we fix $|\tilde{V}_{cs}|=0.975$, as in the SM fit model. We assume lepton-flavour universality of the Wilson coefficients.
@@ 2.5 Hadronic Nuisance Parameters and their Priors
At leading order in $\alpha_{e}$, the leptonic and hadronic matrix elements factorise. Non-factorisable corrections occur only at $\mathcal{O}(\alpha_{e})\simeq 1/137$. Hard virtual correction are included as part of the SM value for the Wilson coefficient $\mathcal{C}^{\ell}_{V,L}$ Sirlin:1980nh while soft real radiation and soft structure-independent corrections are accounted for in the experimental data using the PHOTOS software Golonka:2005pn . Structure-dependent non-factorisable corrections are currently not available for exclusive $c\to s\ell^{+}\nu$ decays.
In our analysis, hadronic matrix elements are described by a substantial number of nuisance parameters. They need to be varied to account for the theoretical uncertainties of the hadronic matrix elements. In the case of leptonic decays, the nuisance parameters are decay constants of the decaying hadrons. In the case of semileptonic decays, the nuisance parameters relate to hadronic form factors, i.e. real-valued functions of $q^{2}=m_{\ell^{+}\nu}^{2}$. The semileptonic nuisance parameters depend on our choice of parametrisation for these form factors. For this analysis, we employ a form factor parametrisation that respects dispersive bounds, which lead to controlled (i.e., parametric) systematic uncertainties in the fit. A summary of the hadronic nuisance parameters pertaining to this analysis is provided below. We refer to appendix A for further details on the definitions of these parameters.
@@
Our analysis uses lattice QCD results by the HPQCD collaboration Parrott:2022rgu for all three form factors, which are provided as parameters of the BCL parametrisation Bourrely:2008za . Since our parametrisation eq. 46 differs from the BCL one, we reconstruct the form factors and their correlations for all three form factors at three different values of the momentum transfer $q^{2}$. We remove one of these points for the form factor $f_{+}^{D\to\bar{K}}$ due to the exact relation between $f_{+}^{D\to\bar{K}}$ and $f_{0}^{D\to\bar{K}}$ at $q^{2}=0$. We further use lattice QCD results by the FNAL/MILC collaboration FermilabLattice:2022gku 222 The HPQCD and FNAL/MILC analyses Parrott:2022rgu ; FermilabLattice:2022gku have a partial overlap in the ensembles used in their calculation. As the number of configurations used in the shared ensembles differs, we estimate that the correlation between these two sets of result can be safely neglected. for the two form factors $f_{+}^{D\to\bar{K}}$ and $f_{0}^{D\to\bar{K}}$, which are also provided in terms of the BCL parameters but with a different truncation order than the HPQCD results. We also reconstruct the form factors and their correlations for both form factors at four different values of the momentum transfer $q^{2}$. Again, the exact form factor relation at $q^{2}=0$ leads us to remove one these $q^{2}$ points. Besides the results by the FNAL/MILC and HPQCD collaborations, results from the ETM collaboration are also available Lubicz:2017syv ; Lubicz:2018rfs . These results provide a total of 8 data points across all three form factors. We comment on the compatibility between the different lattice determinations below.
@@
Figure 1: Prior-predictive distributions for the $D\to\bar{K}$ form factors $f_{+}$, $f_{0}$, and $f_{T}$ as functions of the squared momentum transfer $q^{2}$ at truncation order $K=3$. The 68% probability envelopes for the “nominal” results based are shown as light-blue bands, the envelopes for the ETM results are shown as purple bands; see the discussion in section 2.5. The data points by the ETM Lubicz:2017syv ; Lubicz:2018rfs , FNAL/MILC FermilabLattice:2022gku , and HPQCD Parrott:2022rgu collaborations are overlaid.
@@
A simultaneous fit with our parametrisation to both HPQCD and FNAL/MILC constraints yields acceptable agreement, with a total $\chi^{2}=9.8$ for $4$ degrees of freedom and a $p$ value of about $4\%$, above our a-priori threshold of $3\%$. However, the ETM results are not mutually compatible with the FNAL/MILC and HPQCD results, producing a total $\chi^{2}=63.46$ for $10$ degrees of freedom in a simultaneous fit ($p$ value below $10^{-9}$). We show the data points corresponding to all three lattice QCD results in fig. 1. The tension between ETM on the one hand and FNAL/MILC & HPQCD on the other hand is clearly visible, especially at $q^{2}$ close to its maximum value.333 We also note that the ETM predicted differential rates agree very poorly with the experimental distributions. Comparison plots are provided in the supplementary material EOS-DATA-2024-01 .
@@
In principle, this recipe does not apply to our case, since we aim to combine three multivariate Gaussian distribution. Nevertheless, we adopt the PDG procedure, since there is no standard procedure for our case. We obtain $S^{2}=6.346$. Such a large value further indicates that the ETM results and the FNAL/MILC & HPQCD results are incompatible and warrant further investigations on the lattice side.
This substantial tension between the individual lattice QCD results leads to define our “nominal” scenario. In this scenario, we assume that the ETM results are an outlier and therefore drop them entirely. Hence, we exclusively use the combination of FNAL/MILC and HPQCD results.

### v2

@@ 1 Introduction
The Cabibbo-Kobayashi-Maskawa (CKM) quark mixing matrix is a central parameter in the Standard Model of particle physics (SM). Precision determinations of its elements are key to understanding the origin of (quark) flavour. The CKM element $|V_{cs}|$ is currently the second-most precisely known CKM element. The Particle Data Group (PDG) has determined a reference value for $|V_{cs}|$ with a relative uncertainty of $0.6\%$ PDGCKMReview . In computing this value, the contemporary world averaged of the total branching ratios of $D\to\bar{K}\ell^{+}\nu$ and $D_{s}^{+}\to\ell^{+}\nu$ decays are used in combination with theoretical determinations of the relevant hadronic form factors and decay constants.
In this work, we study exclusive flavour-changing $c\to s\ell^{+}\nu$ processes in a global fit. These processes include $\bar{D}_{s}\to\ell^{+}\nu$, $\bar{D}_{s}^{*}\to\ell^{+}\nu$, $D\to\bar{K}\ell^{+}\nu$, and $\Lambda_{c}\to\Lambda\ell^{+}\nu$. First, we determine if the available data can be simultaneously described by the available theory predictions. This question is not trivial since the hadronic matrix elements are connected within the framework of dispersive bounds Okubo:1973tj ; Boyd:1997kz ; Caprini:1997mu .
We determine $|V_{cs}|$ from the global fit within this setup and discuss the agreement of the result with CKM unitarity. We also probe if determinations of $|V_{cs}|$ from the individual decay processes agree with each other and with the quoted PDG value PDGCKMReview . Finally, we determine how Beyond the Standard Model (BSM) effects are constrained by $c\to s\ell^{+}\nu$ processes by determining the maximally allowed parameter space within the $sc\nu\ell$ sector of the Weak Effective Theory (WET) Aebischer:2017gaw ; Jenkins:2017jig ; Jenkins:2017dyc . Our analysis leverages procedures and techniques first developed in the context of a study of $b\to u\ell^{-}\bar{\nu}$ decays Leljak:2023gna .
@@ 2.1 Theoretical Framework
where $\mathcal{C}^{\ell}_{i}$ are the Wilson coefficients and $\mathcal{O}_{i}$ the local field operators. The normalisation in terms of the Fermi constant $G_{\textrm{F}}$ and an arbitrary constant $\tilde{V}_{cs}$ proves to be convenient later on. The Wilson coefficients encode the dynamics of the full theory, either the SM or any viable BSM theory, above the separation scale $\mu_{c}\simeq 1.275\,\textrm{GeV}$. The matrix elements of the operators encode the dynamics below the separation scale. Here, we assume that there are no right-handed neutrinos of mass less than $\mu_{c}$. Consequently, the basis of operators of mass dimension six reduces to five independent operators. Our choice of basis reads
@@
while all other Wilson coefficients are zero. In the SM, we identify $\tilde{V}_{cs}^{*}$ with the conjugated CKM matrix element $V_{cs}^{*}$. Beyond the SM, the Wilson coefficients provide a low-energy footprint of the genuine BSM dynamics at or above the electroweak scale. These coefficients can then be used to constrain the parameters of a UV-complete BSM model.
@@ 2.2 Approach
Here, $\vec{\vartheta}$ represents the parameters of interest, and $\vec{\nu}$ represents the nuisance parameters. The latter arise exclusively from our description of hadronic matrix elements in the observables. The posterior PDF is proportional to the likelihood function $P(D\,|\,\vec{\vartheta},\vec{\nu},M)$, which encapsulates the experimental and theoretical constraints imposed by the data $D$ under consideration, and to the prior PDF $P_{0}(\vec{\nu},\vec{\vartheta}\,|\,M)$, which accounts for our prior knowledge about the parameters in the model $M$. We discuss the various datasets and fit models entering our analysis in section 2.3 and section 2.4, respectively.
@@
Their ratio, the so-called Bayes factor $K\equiv P(D\,|\,M_{1})/P(D\,|\,M_{2})$, then provides information about the efficiency of the two models in describing the common dataset $D$. Following Jeffreys’ interpretation Jeffreys:1939xee , the model $M_{1}$ is preferred over the model $M_{2}$ if $K>1$. This preference can be characterised either as strong if $10010<K<$ 100 or as decisive if $100<K$. A Bayes factor $103<K<$ 10 is still interpreted to be substantially in favour of $M_{1}$ over $M_{2}$, whereas a Bayes factor of $31<K<$ 3 is “barely worth mentioning”. For $K<1$, the above interpretation favours $M_{2}$ over $M_{1}$ with the replacement $K\to 1/K$.
@@ 2.3 Experimental Data
However, the total decay width of the $D_{s}^{*+}$ is currently unknown from any experiment. Instead, we use the prediction for the decay width determined in Ref. Meng:2024gpd . This value is obtained from the PDG world average of the experimental measurements of the dominant branching ratio $\mathcal{B}(D_{s}^{*+}\to D_{s}^{+}\gamma)$ ParticleDataGroup:2022pth and a lattice QCD prediction for the partial decay width $\Gamma(D_{s}^{*+}\to D_{s}^{+}\gamma)$ Meng:2024gpd .
@@
which is based on results by the Belle Belle:2006idb , BES BES:2004rav , BESIII BESIII:2015tql ; BESIII:2021mfl , and CLEO-c CLEO:2009svp experiments. This average is dominated by the BESIII measurement in Ref. BESIII:2015tql .
@@
based on measurements from BES BES:2004obp , BESIII BESIII:2015jmz ; BESIII:2016hko ; BESIII:2017ylw ; BESIII:2021mfl . The differential decay rate is also available from Ref. BESIII:2017ylw , which we implement as the normalised decay rate $1/\Gamma\,d\Gamma/dq^{2}$. For $D^{+}\to\bar{K}^{0}\mu^{+}\nu$, we use the only available results coming from BESIII BESIII:2016gbw ,
@@
For $\Lambda_{c}\to\Lambda\{e^{+},\mu^{+}\}\nu$ decays, we use measurements of the branching ratios by the BESIII experiment BESIII:2022ysa ; BESIII:2023vfi , yielding
@@
$\mathcal{B}(\Lambda_{c}\to\Lambda\mu^{+}\nu)$ $=(3.48\pm 0.14\pm 0.10)\%\,.$
Although measurements of the differential distributions for the decay chain $\Lambda_{c}\to\Lambda(\to p\pi)e^{+}\nu$ have been undertaken by BESIII BESIII:2022ysa ; BESIII:2023vfi , the data have not been made public. The differential data strongly constrain the BSM parameters space, as discussed in the context of $\Lambda_{b}\to\Lambda_{c}(\to\Lambda\pi)\ell^{-}\bar{\nu}$ Boer:2019zmp ; the analogue of the $\Lambda_{c}$ decays among semileptonic $b\to c\ell^{-}\bar{\nu}$ processes. BESIII also recently measured for the first time the leptonic and hadronic asymmetries of $\Lambda_{c}\to\Lambda\ell\nu$ decays BESIII:2023vfi . However, we cannot use these measurements in our global analysis, because of the absence of publicly available information on their correlations.
@@
We do not use decays to $\eta^{(\prime)}$ because their form factors, while available from light-cone sum rules and lattice QCD Duplancic:2015zna ; Hu:2021zmy ; Bali:2014pva , are not yet as stringently constrained as the ones for the other modes used in this analysis. We also do not use measurements of $D^{+}\to\bar{K}^{*0}\ell^{+}\nu$ nor $D_{s}^{+}\to\phi\ell^{+}\nu$, because their hadronic final states are unstable vector resonances. As such, they may suffer from $S$-wave pollution, which is a problem from the point of view of both the experimental extraction of the observables and the theoretical predictions. Therefore, we do not consider them to be competitive with the processes included above.
@@ 2.4 Parameters of Interest and their Priors
This model has no parameters of interest. We use a fixed normalisation $\tilde{V}_{cs}=0.975$, corresponding to the average determined by the Particle Data Group PDGCKMReview . The left-handed Wilson coefficient $\mathcal{C}^{\ell}_{V,L}$ is set to its SM value in (3), and all other Wilson coefficients are fixed to zero. This fit model serves as the null hypothesis for our model comparisons.
@@
This model has nine parameters of interest, which describe the degrees of freedom for the five complex Wilson coefficients that appear in the effective Hamiltonian eq. 1. The tenth degree of freedom, which can be chosen by imposing $\arg\tilde{V}_{cs}=0=\arg{\mathcal{C}^{\ell}_{V,L}}$, is fixed since the overall phase of the effective Hamiltonian is unobservable. To provide an absolute scale for the magnitude of these Wilson coefficients, we fix $|\tilde{V}_{cs}|=0.975$, as in the SM fit model. We assume lepton-flavour universality of the Wilson coefficients.
@@ 2.5 Hadronic Nuisance Parameters and their Priors
At leading order in $\alpha_{e}$, the leptonic and hadronic matrix elements factorise. Non-factorisable corrections occur only at $\mathcal{O}(\alpha_{e})\simeq 1/137$. Hard virtual corrections are included as part of the SM value for the Wilson coefficient $\mathcal{C}^{\ell}_{V,L}$ Sirlin:1980nh while soft real radiation and soft structure-independent corrections are accounted for in the experimental data using the PHOTOS software Golonka:2005pn . Structure-dependent non-factorisable corrections are currently not available for exclusive $c\to s\ell^{+}\nu$ decays.
In our analysis, hadronic matrix elements are described by a substantial number of nuisance parameters. They need to be varied to account for the theoretical uncertainties of the hadronic matrix elements. In the case of leptonic decays, the nuisance parameters are decay constants of the decaying hadrons. In the case of semileptonic decays, the nuisance parameters relate to hadronic form factors, i.e. real-valued functions of $q^{2}=m_{\ell^{+}\nu}^{2}$. The semileptonic nuisance parameters depend on our choice of parametrisation for these form factors. For this analysis, we employ a form factor parametrisation that respects dispersive bounds, leading to controlled (i.e., parametric) systematic uncertainties in the fit. A summary of the hadronic nuisance parameters pertaining to this analysis is provided below. We refer to appendix A for further details on the definitions of these parameters.
@@
Our analysis uses lattice QCD results by the HPQCD collaboration Parrott:2022rgu for all three form factors, which are provided as parameters of the BCL parametrisation Bourrely:2008za . Since our parametrisation eq. 46 differs from the BCL one, we reconstruct the form factors and their correlations for all three form factors at three different values of the momentum transfer $q^{2}$. We remove one of these points for the form factor $f_{+}^{D\to\bar{K}}$ due to the exact relation between $f_{+}^{D\to\bar{K}}$ and $f_{0}^{D\to\bar{K}}$ at $q^{2}=0$. We further use lattice QCD results by the FNAL/MILC collaboration FermilabLattice:2022gku 222 The HPQCD and FNAL/MILC analyses Parrott:2022rgu ; FermilabLattice:2022gku have a partial overlap in the ensembles used in their calculation. As the number of configurations used in the shared ensembles differs, we estimate that the correlation between these two sets of results can be safely neglected. for the two form factors $f_{+}^{D\to\bar{K}}$ and $f_{0}^{D\to\bar{K}}$, which are also provided in terms of the BCL parameters but with a different truncation order than the HPQCD results. We also reconstruct the form factors and their correlations for both form factors at four different values of the momentum transfer $q^{2}$. Again, the exact form factor relation at $q^{2}=0$ allows us to remove one of these $q^{2}$ points. Besides the results of the FNAL/MILC and HPQCD collaborations, results from the ETM collaboration are also available Lubicz:2017syv ; Lubicz:2018rfs . These results provide a total of 8 data points across all three form factors. We comment on the compatibility between the different lattice determinations below.
@@
Figure 1: Prior-predictive distributions for the $D\to\bar{K}$ form factors $f_{+}$, $f_{0}$, and $f_{T}$ as functions of the squared momentum transfer $q^{2}$ at truncation order $K=3$. The 68% probability envelopes for the “nominal” results are shown as light-blue bands, the envelopes for the ETM results are shown as purple bands; see the discussion in section 2.5. The data points by the ETM Lubicz:2017syv ; Lubicz:2018rfs , FNAL/MILC FermilabLattice:2022gku , and HPQCD Parrott:2022rgu collaborations are overlaid.
@@
A simultaneous fit with our parametrisation to both HPQCD and FNAL/MILC constraints yields acceptable agreement, with a total $\chi^{2}=9.8$ for $4$ degrees of freedom and a $p$ value of about $4\%$, above our a-priori threshold of $3\%$. However, the ETM results are not mutually compatible with the FNAL/MILC and HPQCD results, producing a total $\chi^{2}=63.46$ for $10$ degrees of freedom in a simultaneous fit ($p$ value below $10^{-9}$). We show the data points corresponding to all three lattice QCD results in fig. 1. The tension between ETM on the one hand and FNAL/MILC & HPQCD on the other hand is clearly visible, especially at $q^{2}$ close to its maximum value.333 We also note that the ETM-predicted differential rates agree very poorly with the experimental distributions. Comparison plots are provided in the supplementary material EOS-DATA-2024-01 .
@@
In principle, this recipe does not apply to our case since we aim to combine three multivariate Gaussian distributions. Nevertheless, we adopt the PDG procedure since there is no standard procedure for our case. We obtain $S^{2}=6.346$. Such a large value further indicates that the ETM and FNAL/MILC & HPQCD results are incompatible and warrant further investigations on the lattice side.
This substantial tension between the individual lattice QCD results leads us to defining our “nominal” scenario, in which we assume that the ETM results are an outlier and, therefore, drop them entirely. Hence, we exclusively use the combination of FNAL/MILC and HPQCD results.